# Hybrid A* Path Planning with Multi-Modal Motion Extension for Four-Wheel Steering Mobile Robots

### v2

@@ 1 Introduction
Owing to these multimodal motion capabilities, motion planning for 4WIS systems becomes particularly challenging and critical. Motion planning is generally divided into three levels: global path planning, local path planning and path tracking. Existing studies, however, are largely concentrated on the latter two. For example, Wang et al. combined fuzzy logic with deep reinforcement learning to achieve multimodal local path planning and dynamic decision-making [5]. In addition, various tracking control approaches have been proposed, including backstepping control for symmetric counter-steering [6], sliding mode control for reverse-steering trajectories [7], reinforcement learning integrated with model predictive control (MPC) for counter-steering maneuvers [8], and MPC-based frameworks for stable high-speed trajectory tracking in reverse steering mode [9], among others. These efforts have substantially advanced the local planning and tracking capabilities of 4WIS robots.
Despite these advances, research on long-range global planning that fully exploits the multimodal capabilities of 4WIS systems remains limited. Classical planners such as RRT* [10] and A* [11] typically neglect kinematic constraints; optimization-based methods suffer from slow convergence and high computational burden in large-scale scenarios; and learning-based approaches are constrained by memory horizons, which limit their ability to generalize beyond local contexts.
Hybrid A* [12], which combines graph search with sampling-based strategies, has emerged as an effective approach to balance computational efficiency and scalability while respecting robot kinematics. However, standard Hybrid A* does not explicitly account for the multimodal motion mechanisms inherent to 4WIS systems, leaving much of their maneuvering potential underutilized. In our previous work [13], we conducted a preliminary extension of Hybrid A* by incorporating selected 4WIS motion modes into the motion primitive set, thereby enabling rudimentary multimodal planning. Nevertheless, this approach remained exploratory: the heuristic function was simplified, and issues such as efficient mode switching and integration of deeper kinematic constraints were not systematically addressed.
To overcome these limitations, this paper proposes a systematic multimodal motion planning framework that extends Hybrid A* by redesigning the heuristic function, optimizing motion primitive generation, and introducing an explicit mode-switching strategy. In particular, the enhanced heuristic accelerates global search by better capturing the kinematic characteristics of four-wheel independent steering robots, while the optimized motion primitives improve path smoothness and adaptability across different maneuvering modes. Furthermore, the proposed mode-switching mechanism explicitly coordinates transitions between steering modes, ensuring both feasibility and efficiency.
@@ 2.1 Global Path Planning Problem Formulation
$\begin{cases}\dot{x}=\frac{1}{4}\sum_{i=1}^{4}v\cos(\eta_{i}+\theta)\\ \dot{y}=\frac{1}{4}\sum_{i=1}^{4}v\sin(\eta_{i}+\theta)\\ \dot{\theta}=\sum_{i=1}^{4}\frac{v(-y^{r}_{i}\cos\eta_{i}+x^{r}_{i}\sin\eta_{i})}{4{x^{r}_{i}}^{2}+4{y^{r}_{i}}^{2}},\end{cases}$ (4)
where $v_{i}$ and $\eta_{i}$ denote the linear velocity and steering angle of the $i$-th wheel. The subscripts 1, 2, 3, and 4 correspond to the front-left, front-right, rear-left, and rear-right wheels, respectively.
@@
Ackermann steering achieves smooth cornering by directing all wheels toward a common instantaneous center of rotation. For given control inputs $\{v,\eta\}$, the relationships between wheel steering angles and velocities are:
$\begin{cases}\cot\eta_{1}=-\cot\eta_{3}=\cot\eta-W/L\\ \cot\eta_{2}=-\cot\eta_{4}=\cot\eta+W/L\\ v_{1}=v_{3}=v\tan\eta/\sin\eta_{1}\\ v_{2}=v_{4}=v\tan\eta/\sin\eta_{2},\end{cases}$ (5)
where $v$ denotes the robot’s longitudinal velocity, $\eta$ represents the desired steering angle, $\eta_{i}$ is the steering angle of the $i$-th wheel, and $v_{i}$ is the linear velocity of the $i$-th wheel.
Lateral steering preserves the same same geometric constraint principles as the Ackermann mode, but the wheels are reoriented to be perpendicular to the vehicle’s longitudinal axis, allowing the robot to generate pure sideways motion. When the lateral direction is regarded as the effective forward direction, the underlying steering-angle computation becomes structurally identical to that of the Ackermann formulation, with $L$ and $W$ exchanging roles in the geometric constraint. In parallel movement mode, all wheels maintain identical steering angles and velocities, so that the vehicle undergoes pure translation in an arbitrary direction without changing its orientation.
@@ 3.2 Multi-Modal State Representation
In the expanded four-dimensional state space, state expansion process includes two operations: intra-modal expansion and inter-modal expansion. Intra-modal expansion maintains the current motion mode unchanged and generates successor states in the three-dimensional pose space according to the corresponding kinematic model; inter-modal expansion performs mode switching at the current position, producing successor states with different motion modes but the same pose. To effectively manage state space complexity, the algorithm employs heuristic modal selection strategies, choosing the most effective motion mode based on the relative direction between the current position and target, and minimizing unnecessary modal switches while maintaining trajectory continuity under constraint satisfaction. The sampling strategies associated with each motion mode are illustrated in Fig. 2.
@@ 3.3 Multi-Modal Reeds-Shepp Curves
To accommodate the multi-modal motion characteristics of four-wheel independent steering systems, this research extends the traditional RS curve calculation method. The generation of RS curves depends on key parameters such as minimum turning radius, and different motion modes have different kinematic characteristics, thus requiring separate design of corresponding curve calculation methods.
@@
$\kappa^{(1)}=\frac{2\tan(\delta_{\max})}{L},$ (8)
where $\delta_{\max}$ is the maximum front wheel steering angle.
@@ 3.4 Cost and Heuristic Functions
The cost function is extended by introducing a mode switching penalty into the traditional $g(n)$. The calculation of mode switching cost directly utilizes the aforementioned reference velocity and sampling time mechanism: $C_{\text{switch}}=v_{\text{ref}}\times t_{\text{switch}}$, where $t_{\text{switch}}$ is the preset switching time. This time-velocity based quantification method enables mode switching costs to be compared with path length in the same dimension. The extended accumulated cost function is expressed as:
$+C_{\text{steer-change}}\cdot|\Delta\delta|$
@@ 3.5 Terminal Connection Strategy
19:     Generate inter-modal successor (mode switch)
@@ 4.1 Experimental Setup
Max Steer Angle ($\delta_{\max}$) 30° Max Steer Rate ($\dot{\delta}_{\max}$) 360°/s
Reference Velocity ($v_{\text{ref}}$) 1.0 m/s Sampling Time ($\Delta t$) 0.4 s
Mode Switch Time ($t_{\text{switch}}$) 1.0 s Wheel Radius ($r_{w}$) 0.13 m
@@ 4.2 Performance Comparison
In the maze environment, the comparative results are shown in Fig. 4. Due to the presence of narrow corridors and sharp turns, the baseline Hybrid A* relying solely on Ackermann steering often produces longer and more tortuous paths. In contrast, the proposed multimodal Hybrid A* can flexibly switch motion modes according to the scenario. For example, in Env1-S1 and Env1-S2, the lateral mode enables the robot to directly traverse narrow passages, resulting in shorter and smoother paths. In the more challenging Env1-S3, the planner seamlessly integrates Ackermann and lateral maneuvers, effectively reducing steering costs while maintaining path feasibility.
In the parking environment, the comparison is illustrated in Fig. 5. For side-parking tasks (Env2-S1 and Env2-S2), the baseline Hybrid A* frequently requires multiple adjustments and backward maneuvers, whereas the multimodal planner leverages lateral and translational movements to achieve more efficient parking operations. In the long-distance slot-changing scenario (Env2-S3), the extended planner further demonstrates the advantage of motion mode combinations, significantly reducing unnecessary reversing and large steering maneuvers. This capability highlights its superior mobility and operational efficiency in structured environments.
Table 3 summarizes the quantitative differences between the two approaches. Across all test cases, the proposed method consistently outperforms the baseline in terms of both path length and path cost, with the most significant improvements observed in cost reduction. This is attributed to the cost function explicitly penalizing excessive steering and reversing, while the multimodal switching strategy effectively mitigates these penalties. Overall, the results confirm that the proposed framework not only enhances geometric efficiency but also demonstrates strong generality and robustness across diverse tasks, providing practical potential for 4WIS robots operating in complex environments.
Table 3: Performance Comparison between Hybrid A* and Extended Multi-modal Hybrid A* in Various Scene
Scenario Algorithm Path Length (m) Path Cost
Env1-S1 Hybrid A* 18.27 36.79
Multi-moda Hybrid A* 15.95 20.28
Env1-S2 Hybrid A* 15.27 25.85
Multi-modal Hybrid A* 14.80 19.70
Env1-S3 Hybrid A* 46.40 73.13
Multi-modal Hybrid A* 45.55 58.35
Env2-S1 Hybrid A* 3.55 12.99
Multi-modal Hybrid A* 3.33 10.82
Env2-S2 Hybrid A* 7.25 23.35
Multi-modal Hybrid A* 7.15 11.84
Env2-S3 Hybrid A* 17.92 34.33
Multi-modal Hybrid A* 17.54 32.91
@@
To further validate the feasibility of the proposed framework in practical applications, we conducted preliminary experiments on a physical 4WIS platform as shown in Fig. 6. MPC tracking controllers were constructed for different modes, and tracking tests were performed on six multi-modal trajectories planned in simulation scenarios. The selected test trajectories cover most typical maneuvering actions and mode switching scenarios, demonstrating good representativeness. Under a reference velocity of 1 m/s, the maximum lateral error was controlled within 4 cm and the maximum longitudinal error was kept below 6 cm, including errors at switching points, which validates the feasibility and effectiveness of the planned trajectories in real systems.

### v3

@@ 1 Introduction
Owing to these multi-modal motion capabilities, motion planning for 4WIS systems becomes particularly challenging and critical. Motion planning is generally divided into three levels: global path planning, local path planning and path tracking. Existing studies, however, are largely concentrated on the latter two. For example, Wang et al. combined fuzzy logic with deep reinforcement learning to achieve multi-modal local path planning and dynamic decision-making [5]. In addition, various tracking control approaches have been proposed, including backstepping control for symmetric counter-steering [6], sliding mode control for reverse-steering trajectories [7], reinforcement learning integrated with model predictive control (MPC) for counter-steering maneuvers [8], and MPC-based frameworks for stable high-speed trajectory tracking in reverse steering mode [9], among others. These efforts have substantially advanced the local planning and tracking capabilities of 4WIS robots.
Despite these advances, research on long-range global planning that fully exploits the multi-modal capabilities of 4WIS systems remains limited. Classical planners such as RRT* [10] and A* [11] typically neglect kinematic constraints; optimization-based methods suffer from slow convergence and high computational burden in large-scale scenarios; and learning-based approaches are constrained by memory horizons, which limit their ability to generalize beyond local contexts.
Hybrid A* [12], which combines graph search with sampling-based strategies, has emerged as an effective approach to balance computational efficiency and scalability while respecting robot kinematics. However, standard Hybrid A* does not explicitly account for the multi-modal motion mechanisms inherent to 4WIS systems, leaving much of their maneuvering potential underutilized. In our previous work [13], we conducted a preliminary extension of Hybrid A* by incorporating selected 4WIS motion modes into the motion primitive set, thereby enabling rudimentary multi-modal planning. Nevertheless, this approach remained exploratory: the heuristic function was simplified, and issues such as efficient mode switching and integration of deeper kinematic constraints were not systematically addressed.
To overcome these limitations, this paper proposes a systematic multi-modal motion planning framework that extends Hybrid A* by redesigning the heuristic function, optimizing motion primitive generation, and introducing an explicit mode-switching strategy. In particular, the enhanced heuristic accelerates global search by better capturing the kinematic characteristics of four-wheel independent steering robots, while the optimized motion primitives improve path smoothness and adaptability across different maneuvering modes. Furthermore, the proposed mode-switching mechanism explicitly coordinates transitions between steering modes, ensuring both feasibility and efficiency.
@@ 2.1 Global Path Planning Problem Formulation
$\begin{cases}\dot{x}=\frac{1}{4}\sum_{i=1}^{4}v\cos(\phi_{i}+\theta)\\ \dot{y}=\frac{1}{4}\sum_{i=1}^{4}v\sin(\phi_{i}+\theta)\\ \dot{\theta}=\sum_{i=1}^{4}\frac{v(-y^{r}_{i}\cos\phi_{i}+x^{r}_{i}\sin\phi_{i})}{4{x^{r}_{i}}^{2}+4{y^{r}_{i}}^{2}},\end{cases}$ (4)
where $v_{i}$ and $\phi_{i}$ denote the linear velocity and steering angle of the $i$-th wheel. The subscripts 1, 2, 3, and 4 correspond to the front-left, front-right, rear-left, and rear-right wheels, respectively.
@@
Ackermann steering achieves smooth cornering by directing all wheels toward a common instantaneous center of rotation. For given control inputs $\{v,\phi\}$, the relationships between wheel steering angles and velocities are:
$\begin{cases}\cot\phi_{1}=-\cot\phi_{3}=\cot\phi-W/L\\ \cot\phi_{2}=-\cot\phi_{4}=\cot\phi+W/L\\ v_{1}=v_{3}=v\tan\phi/\sin\phi_{1}\\ v_{2}=v_{4}=v\tan\phi/\sin\phi_{2},\end{cases}$ (5)
where $v$ denotes the robot’s longitudinal velocity, $\phi$ represents the desired steering angle, $\phi_{i}$ is the steering angle of the $i$-th wheel, and $v_{i}$ is the linear velocity of the $i$-th wheel.
Lateral steering preserves the same same geometric constraint principles as the Ackermann mode, but the wheels are reoriented to be perpendicular to the robot’s longitudinal axis, allowing the robot to generate pure sideways motion. When the lateral direction is regarded as the effective forward direction, the underlying steering-angle computation becomes structurally identical to that of the Ackermann formulation, with $L$ and $W$ exchanging roles in the geometric constraint. In parallel movement mode, all wheels maintain identical steering angles and velocities, so that the robot undergoes pure translation in an arbitrary direction without changing its orientation.
@@ 3.2 Multi-Modal State Representation
In the expanded four-dimensional state space, state expansion process includes two operations: intra-modal expansion and inter-modal expansion. Intra-modal expansion maintains the current motion mode unchanged and generates successor states in the three-dimensional pose space according to the corresponding kinematic model; inter-modal expansion performs mode switching at the current position, producing successor states with different motion modes but the same pose. To effectively manage state space complexity, the algorithm employs heuristic modal selection strategies, choosing the most effective motion mode based on the relative state between the current position and target, and minimizing unnecessary modal switches while maintaining trajectory continuity under constraint satisfaction. The sampling strategies associated with each motion mode are illustrated in Fig. 2.
@@ 3.3 Multi-Modal Reeds-Shepp Curves
To accommodate the multi-modal motion characteristics of four-wheel independent steering systems, this research extends the traditional RS curve calculation method. The generation of RS curves depends on key parameters such as maximum curvature, and different motion modes have different kinematic characteristics, thus requiring separate design of corresponding curve calculation methods.
@@
$\kappa^{(1)}=2\tan(\phi_{\max})/{L},$ (8)
where $\phi_{\max}$ is the maximum wheel steering angle.
@@ 3.4 Cost and Heuristic Functions
The cost function is extended by introducing a mode switching penalty into the traditional $g(n)$. The calculation of the mode switching cost leverages the aforementioned reference velocity and sampling time mechanism and explicitly accounts for the required full stop and subsequent re-acceleration during a mode transition:
$C_{\text{switch}}=v_{\text{ref}}\times t_{\text{switch}}+v_{\text{ref}}^{2}/a_{max},$
where $t_{\text{switch}}$ denotes the preset switching time and $a$ represents the robot acceleration. For mode switching occurring at the initial node, the system starts from rest and only undergoes the acceleration phase. As a result, the acceleration-related component of the switching cost is halved, since no deceleration process is required. This time-velocity based quantification method enables mode switching costs to be compared with path length in the same dimension. The extended accumulated cost function is expressed as:
$g(n)=g_{\text{prev}}+\Delta s+C_{\text{reverse}}\cdot\mathbb{I}_{\{d=-1\}}+C_{\text{steer}}\cdot|\phi|$
$+C_{\text{steer-change}}\cdot|\Delta\phi|$
@@
19:     Generate inter-modal successors (mode switch)
@@ 4.1 Experimental Setup
Max Steer Angle ($\phi_{\max}$) 30° Max Steer Rate ($\dot{\phi}_{\max}$) 360°/s
Reference Velocity ($v_{\text{ref}}$) 1.0 m/s Sampling Time ($\Delta t$) 0.2 s
Mode Switch Time ($t_{\text{switch}}$) 0.5 s Max Acceleration ($a_{\max}$) 2 $\text{m/s}^{2}$
@@ 4.2 Performance Comparison
In the maze environment, the comparative results are shown in Fig. 4. Due to the presence of narrow corridors and sharp turns, the baseline Hybrid A* relying solely on Ackermann steering often produces longer and more tortuous paths. In contrast, the proposed multi-modal Hybrid A* can flexibly switch motion modes according to the scenario. For example, in Env1-S1 and Env1-S2, the lateral mode enables the robot to directly traverse narrow passages, resulting in shorter and smoother paths. In the more challenging Env1-S3, the planner seamlessly integrates Ackermann and lateral maneuvers, effectively reducing steering costs while maintaining path feasibility.
In the parking environment, the comparison is illustrated in Fig. 5. For side-parking and exit tasks (Env2-S1 and Env2-S2), the baseline Hybrid A* frequently requires multiple adjustments and backward maneuvers, whereas the multi-modal planner leverages lateral and translational movements to achieve more efficient parking operations. In the long-distance slot-changing scenario (Env2-S3), the extended planner further demonstrates the advantage of motion mode combinations, significantly reducing unnecessary reversing and large steering maneuvers. This capability highlights its superior mobility and operational efficiency in structured environments.
Table 3 summarizes the quantitative differences between the two approaches. Across all test cases, the proposed method consistently outperforms the baseline in terms of both path length and path cost, with the most significant improvements observed in cost reduction. This is attributed to the cost function explicitly penalizing excessive steering and reversing, while the multi-modal switching strategy effectively mitigates these penalties. Overall, the results confirm that the proposed framework not only enhances geometric efficiency but also demonstrates strong generality and robustness across diverse tasks, providing practical potential for 4WIS robots operating in complex environments.
Table 3: Performance Comparison between Hybrid A* and Multi-modal Hybrid A*
Scenario Hybrid A* Multi-modal Hybrid A*
Path Length Path Cost Path Length Path Cost
Env1-S1 18.49 34.75 15.88 20.03
Env1-S2 14.89 25.89 14.80 19.44
Env1-S3 48.56 59.78 45.35 58.08
Env2-S1 12.69 18.48 12.27 15.32
Env2-S2 6.73 16.04 6.44 9.87
Env2-S3 16.61 34.76 16.49 33.69
@@
To further validate the feasibility of the proposed framework in practical applications, we conducted preliminary experiments on a physical 4WIS platform as shown in Fig. 6. DRL tracking controllers [14] were constructed for different modes, and tracking tests were performed on six multi-modal trajectories planned in simulation scenarios. The selected test trajectories cover most typical maneuvering actions and mode switching scenarios, demonstrating good representativeness. Under a reference velocity of 1 m/s, the average lateral error was controlled within 3 cm and the average longitudinal error was kept below 5 cm, including errors at switching points, which validates the feasibility and effectiveness of the planned trajectories in real systems.